OpenStudy (goldrush18):

given the series 54+18+6+...determine the: i. general term ii. sum to infinity

OpenStudy (anonymous):

this is a GP ,,, u just need to find common ratio

OpenStudy (goldrush18):

and wat is the common ratio

OpenStudy (goldrush18):

dont knw how to do it

OpenStudy (anonymous):

in GP each term after the first is found by multiplying the previous one by a fixed non-zero number called the common ratio

OpenStudy (anonymous):

for example here\[18=54\times\frac{1}{3}\]\[6=18\times\frac{1}{3}\]...

OpenStudy (anonymous):

so the common ratio is \[r=\frac{1}{3}\]

OpenStudy (goldrush18):

wat is the sum of infinity?

OpenStudy (goldrush18):

ok thanx